Question 10 (1 point)

Suppose that a researcher studying the weight of female college athletes wants to predict the weights based on height, measured in inches, and the percentage of body fat of an athlete. The researcher calculates the regression equation as (weight) = 4.582*(height) + 0.953*(percent body fat) - 80.689. If a female athlete is 68 inches tall and has a 18 percentage of body fat, what is her expected weight?

Question 10 options:

1)

We do not know the observations in the data set, so we cannot answer that question.

2)

328.73

3)

66.591

4)

409.419

5)

248.041
Answer #1

Solution:

Given: the regression equation is:

(weight) = 4.582*(height) + 0.953*(percent body fat) - 80.689.

We have to estimate the expected weight of female athlete who is 68 inches tall and has a 18 percentage of body fat.

Thus

(weight) = 4.582*(height) + 0.953*(percent body fat) - 80.689.

(weight) = 4.582*(68) + 0.953*(18) - 80.689.

(weight) =311.576 +17.154 - 80.689.

(weight) = 248.041